OpenStudy (anonymous):

ames creates the table shown to represent the function ƒ(x) = x2 + 8x + 12. Determine the zero(s) of the function.

OpenStudy (anonymous):

x f(x) -8 12 -7 5 -6 0 -5 -3 -4 -4 -3- -3 -2 0 -1 5

sam (.sam.):

Zeros are when f(x)=0 So ƒ(x) = x2 + 8x + 12=0 x2 + 8x + 12=0 Factor it